OpenStudy (anonymous):

Determine the rate law, including the values of the orders and rate law constant, for the following reaction using the experimental data provided.

OpenStudy (anonymous):

Trial [Q] [X] Rate 1 0.12 M 0.10 M 1.5 × 10-3 M/min 2 0.24 M 0.10 M 3.0 × 10-3 M/min 3 0.12 M 0.20 M 12.0 × 10-3 M/min

OpenStudy (anonymous):

do you know how to relate the reaction rate to the concentrations of the reactants?

OpenStudy (jfraser):

to find the order of each reactant, find the ratio of the rates when the concentrations are changed. I'm assuming the reaction is:\[Q + X \rightarrow QX\] so the starting concentrations of Q and X are given, as well as the initial rate

OpenStudy (jfraser):

the equation for a rate law is:\[Rate = k*[Q]^m*[X]^n\]the value of k I don't know, and the values of m and n ARE NOT THE SAME AS THE BALANCED COEFFICIENTS (important to remember that)

OpenStudy (jfraser):

I'm going to divide R2 by R1 and show you what I get

OpenStudy (jfraser):

\[\frac{R_2}{R_1} = \frac{k*[0.24]^m*[0.10]^n}{k*[0.12]^m*[0.10]^n}\]

OpenStudy (anonymous):

k[q][x]^3

OpenStudy (jfraser):

no

OpenStudy (anonymous):

oh ok

OpenStudy (jfraser):

the exponents for the rate law have to be determined using data, so the powers of Q and X are unknown (so far). that's what i'm going to solve for

OpenStudy (anonymous):

ohhh ok

OpenStudy (jfraser):

in the equation i posted, the RATIO of rates is what's important

OpenStudy (jfraser):

\[\frac{R_2}{R_1} = \frac{k*[0.24]^m*[0.10]^n}{k*[0.12]^m*[0.10]^n} = \frac{3.0*10^{-3}}{1.5*10^{-3}} = 2\]

OpenStudy (jfraser):

see what happens to the K's? See what happens to the value of n?

OpenStudy (anonymous):

ther whole #s

OpenStudy (jfraser):

they cancel

OpenStudy (anonymous):

the ns cancel?

OpenStudy (jfraser):

the n's cancel because the concentrations of X are the same. When you divide 0.1 by 0.1, you get 1. You raise 1 to any power and you get 1, so the entire term of X^n cancels

OpenStudy (anonymous):

ok so how do you find the final answer?

OpenStudy (jfraser):

I don't have the final answer yet

OpenStudy (jfraser):

you have to set it up so that all terms cancel except one at a time

OpenStudy (jfraser):

\[\frac{R_2}{R_1} = \frac{k*[0.24]^m*[0.10]^n}{k*[0.12]^m*[0.10]^n} = \frac{3.0*10^{-3}}{1.5*10^{-3}} = 2\]so\[2 = \frac{[0.24]^n}{[0.12]^n} = 2^n\]

OpenStudy (anonymous):

okkkk

OpenStudy (jfraser):

the K's cancel because they're equal. the n's cancel because the concentrations of X are the same, so it's 1^n which always equals 1

OpenStudy (jfraser):

the only terms left over simplify into 2^n = 2. There's only 1 number that makes that true.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do you have the final answer?

OpenStudy (jfraser):

what's the only possible answer for n, in this case?

OpenStudy (anonymous):

n=2

OpenStudy (jfraser):

no\[2^n = 2\] n must equal 1

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

2^n=2

OpenStudy (anonymous):

?

OpenStudy (jfraser):

2^n = 2, so n must equal 1

OpenStudy (jfraser):

do the same thing with R3/R1, to solve for the value of the other power

OpenStudy (anonymous):

divide by 2 on the left

OpenStudy (jfraser):

no. write out the whole equations and see which parts cancel by taking R3/R1

OpenStudy (anonymous):

can you show me?

OpenStudy (jfraser):

do the same thing i did above doing R2/R1, but use the data from R3 on top instead of R2

OpenStudy (anonymous):

\[r3=k*12*20\]

OpenStudy (jfraser):

that's R3, now what's R3/R1? set it up like I did above

OpenStudy (anonymous):

k*12*10

OpenStudy (jfraser):

that's R1, now divide R3/R1\[\frac {R_3}{R_1} = \]

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

okk,then what?

OpenStudy (jfraser):

simplify and solve for n

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

1

OpenStudy (jfraser):

sorry, made a typo. it should be: \[\frac{R_3}{R_1} = \frac{k*[0.12]^m*[0.20]^n}{k*[0.12]^m*[0.10]^n} = \frac{12*10^{-3}}{1.5*10^{-3}}\] the k's cancel, the m's cancel. all you're left with is\[\frac{R_3}{R_1} = \frac{[0.20]^n}{[0.10]^n} = \frac{12*10^{-3}}{1.5*10^{-3}}\]

OpenStudy (anonymous):

how do we get past this part?

OpenStudy (jfraser):

simplify and solve for n. algebra

OpenStudy (anonymous):

can you help?

OpenStudy (jfraser):

\[\frac{R_3}{R_1} = (\frac{[0.20]}{[0.10]})^n = \frac{12*10^{-3}}{1.5*10^{-3}} = 8\] \[2^n = 8\]

OpenStudy (anonymous):

thats not the final tho

OpenStudy (jfraser):

that's true. what's the only value of n that makes that equation true?

OpenStudy (anonymous):

umm? lol

OpenStudy (anonymous):

4?

OpenStudy (jfraser):

2x4 is 8, but 2 RAISED to the 4th is not 8

OpenStudy (anonymous):

ok so? im not sure :/

OpenStudy (jfraser):

\[2^1= 2\]\[2^2 = 2 * 2\]\[2^3 = 2*2*2\]\[2^4 = 2*2*2*2\]

OpenStudy (anonymous):

oh okkk that mases sence

OpenStudy (jfraser):

so what value of 2^n = 8?

OpenStudy (anonymous):

2^4=8

OpenStudy (anonymous):

?

OpenStudy (jfraser):

2*2*2*2 = 8?

OpenStudy (anonymous):

3

OpenStudy (jfraser):

n = 3

OpenStudy (anonymous):

oh okk so what would be the fnal answer?

OpenStudy (jfraser):

What's the value of k? the powers of Q and X are now solved for, so find the value of k

OpenStudy (anonymous):

okk so whats next?

OpenStudy (jfraser):

plug in a set of values and solve for k. you've got 3 trials to pick from

OpenStudy (anonymous):

how do we do yhat?

OpenStudy (jfraser):

pick one of the trials. plug in numbers into the equation for rate. solve for k

OpenStudy (anonymous):

can you set one up for me?

OpenStudy (jfraser):

i think you should try setting it up first

OpenStudy (australopithecus):

Here is a chapter on kinetics that should be of interest to you, Jfraser seems to be able to help you at the moment im in lecture so I cant really be of help to you.

OpenStudy (australopithecus):

He has pretty much given you the answer just take a break and come back to it :) Im sure you can get it.

OpenStudy (anonymous):

ive been trying lol

OpenStudy (anonymous):

r=2*12*10 ?

OpenStudy (jfraser):

\[R = k*[Q]^1*[X]^3\]\[1.5*10^{-3} = k*[0.12]^1*[0.10]^3\]

OpenStudy (anonymous):

how do i know if its right

OpenStudy (jfraser):

because i've done all the work. it's right

OpenStudy (jfraser):

solve for k

OpenStudy (anonymous):

10^-3 = -1000?

OpenStudy (jfraser):

no. 10^-3 is 0.001

OpenStudy (anonymous):

0.0015

OpenStudy (anonymous):

0.0015=k*.00012

OpenStudy (anonymous):

k=.08

OpenStudy (jfraser):

no. you flipped it upside down

OpenStudy (anonymous):

? how lol

OpenStudy (anonymous):

oh i see the right side is positive 10^3

OpenStudy (anonymous):

-.00012

OpenStudy (anonymous):

-0.08 ?

OpenStudy (jfraser):

it's upside down, not negative. If\[0.0015 = k*0.00012\]then\[k = \frac{0.0015}{0.00012}\] you flipped the fraction upside down

OpenStudy (anonymous):

so it a whole #

OpenStudy (anonymous):

k=12.5

OpenStudy (jfraser):

it's not guaranteed to be a whole number, just larger than 1 k =12.5